# The Eye as a Window to Systemic Health: A Survey of Retinal Imaging from Classical Techniques to Oculomics

### v1

@@ II-B8 Overview on the Evolution of AI
In a nutshell, the emergence of AI has reshaped the dynamics of almost every field, specifically healthcare. The short description, limitations, evolution, and relevance to medicine are shown in Figure 3, and some related stuff is discussed here. In 1956, the emergence of AI took its main start from the Dartmouth College conference, which mainly consisted of symbolic AI, in other words, a rule-based system. However, due to the computational limit, the transition of the rule-based system into the classical ML occurred from 1966 to 1970. The wave of backpropagation was another milestone in 1980; despite the emergence of this technique, AI was not getting full attention due to the funding constraints, along with the expected results [59]. The era 1980-1990 is considered a winter era for AI, and again, the domain is shifted to statistical and classical ML. The Chronology of LENET Architecture, which spans from 1990 to 2012, laid the foundation for the emergence of the DL model and paved a new horizon for the future of AI [60, 61]. The DL model is a seismic change that is now non-stop in the current era, and day-to-day, new algorithms and models are emerging. The transition of AI from a theoretical to a practical tool has changed the dynamics of the research, academia and industrial world. The capabilities of AI now extend from the basic models to the complex, whether supervised or unsupervised.
@@
III Modern Techniques: The Rise of Oculomics
As mentioned earlier, the eye disease is retained in the retina. The same is also connected and affects the important parts of the body, such as the brain, heart, kidneys, etc. This type of disease is called systemic disease, a condition that affects the entire body or several organs, not just one part. The linking of the eyes with systemic diseases is named oculomics [1, 4]. The potential behind the oculomics expanded further by leveraging the existing retinal imaging research with the advancement of multiple data modalities. This can be gained by revealing the medical modalities. The modalities used so far in the field of ophthalmology are OCT, fluorescence angiography (FA), adaptive optics (AO), and FP for revealing the structural and functional changes in the ocular system [2, 3]. These modalities help provide enormous retinal imaging data and are widely used in the real world of medical vision. This imaging data emerged as the dataset used to mitigate the burden of eye diseases with respect to monitoring, therapeutic, prognostic and diagnosis, respectively [2]. The datasets are used in different case studies individually and combined for monitoring and identifying a variety of eye ailments like DR, GL, ME, and cataracts etc. The different retinal datasets are available publically to detect and classify eye diseases on time. These datasets have been very helpful so far and are used for different objectives of classical ocular diseases.
III-A Retinal Imaging Techniques
OCT, FP, FA, and AO are the primary medical imaging modalities used to screen for visual disorders. OCT and FP, among the medical imaging modalities, play a pivotal role in screening visual disorders. Their non-invasive nature and detailed interpretation make them stand out. OCT, in particular, has become the gold standard for visualising retinal structures at micrometre resolution. The ability to provide cross-sectional images allows clinicians to analyze retinal layers for conditions such as macular edema, glaucoma, and AMD. The extended version of OCT, OCTA, captures the underlying mechanism of the blood flow of choroidal and retinal vasculature, providing valuable insights into retinal neovascularization and ischemic conditions [62].
For the past three decades of epidemiological studies in developed countries, fundus photography has been used as a standard for processing retinal imaging subjects to quantify the retinal vascular geometric parameters, including vessel tortuosity, fractal dimension (FD), width, diameter, and focal narrowing. The intuition behind the fundus is that the retina is the only part of the central nervous system (CNS) which can be seen through the ophthalmoscope. When the ophthalmologist looks through the ophthalmoscope into the eyeball, they are able to visualise the posterior part of the retina and the structure around it. This part of the retina, which is visible to the telescope, is called the ocular fundoscopy. FP remains one of the most widely used tools in retinal imaging. It captures detailed 2D retina images, including the optic disc, macula, and retinal vasculature. It is invaluable for screening conditions like DR, hypertensive retinopathy, and retinal vein occlusion. However, fundus photography’s inability to provide depth information limits its application in detecting subtle changes in retinal layers. Similarly, Fluorescein angiography uses a fluorescent dye to highlight retinal and choroidal vasculature. It has been a critical tool for diagnosing retinal vein occlusion, AMD, and diabetic ME. Despite its diagnostic utility, the invasive nature of dye injection can pose risks to patients. Another non-invasive technique is AO. AO imaging corrects optical distortions to achieve cellular-level resolution. This technique allows for the visualisation of individual photoreceptors and capillary networks, providing insights into microvascular changes and cellular damage in diseases like hypertension and DR stages. Integrating any of the above approaches is also helpful in some cases [63] By combining imaging modalities like OCT, FP, and FA, we can comprehensively understand retinal structure and function. This multimodal imaging approach is increasingly being used in clinical practice, enhancing our ability to diagnose complex conditions with greater precision. Studies have shown that this technique significantly improves the efficiency of our diagnostic models compared to using a single imaging modality, providing reassurance about the progress we are making in the field [64].
III-B Publicly Available Retinal Datasets
In the present day, the data is used as a tool for competing, from small tech companies to giants and from developing to developed countries. Data is valuable and works as a gold mine in the epoch of GEN-AI. In this context, the available data is getting more attention if found in large amounts, specifically in the field of the medical sciences. This data can emerge as a dataset which can be used further in the real world of health research and drive innovations and discoveries [19]. Furthermore, gold-standard imaging techniques provide the opportunity for the creation of a dataset used as a benchmark for the validation and testing of real clinical trials in situ and in vivo. The advancement of retinal imaging techniques has created numerous publicly available datasets, which have become indispensable resources for researchers and physicians. These datasets are widely used for lesion detection, disease classification, and vessel segmentation [65]. For example, datasets such as STARE [66],DRIVE [67],DIARETDB1 [68] ,CHASE DB1, [69] ,HRF [70], E-ophtha [71] , MESSIDOR, MESSIDOR-2 [72] ,KAGGLE-EyePACS [73] ,IDRiD [74], KAGGLE-APTOS [75] , and DDR [76] have been instrumental in developing algorithms for detecting and classifying diabetic retinopathy, segmenting retinal vessels, and identifying other ocular pathologies.
In this survey, we therefore transition from the dataset overview into a focused discussion of key localized diseases—beginning with glaucoma, where structural deterioration of the optic nerve is a primary concern; followed by age-related macular degeneration, which involves degeneration of the macula; and finally diabetic retinopathy, where vascular abnormalities and lesion characteristics dominate. Each disease presents unique imaging biomarkers and machine learning challenges, necessitating tailored approaches in algorithm design and data representation.
Figure 4: An annotated fundus image reproduced from Li et al. [65], showing pathological markers including MAs, HMs, EXs, and anatomical regions like the fovea and optic disc.
Glaucoma: Glaucoma is a long-term and gradually worsening condition of the optic nerve that results in permanent loss of vision and stands as one of the primary causes of blindness worldwide. It affects an estimated 76 million people as of 2020 and is projected to impact over 111.8 million by 2040 [77]. Elevated intraocular pressure is often linked to the disease, but many patients also experience normal-tension glaucoma, illustrating the condition’s complex origins. The disease affects both the structural and functional health of the eye, especially through the degeneration of retinal ganglion cells (GC) and harm to the optic nerve head. Clinically, a key biomarker for diagnosing glaucoma is the cup-to-disc ratio (CDR), which quantifies the size of the optic cup relative to the optic disc and reflects the cupping of the optic nerve [78, 79].
Two primary imaging modalities, FP and OCT, are widely used to detect and monitor glaucomatous changes. OCT provides high-resolution cross-sectional views of the retinal nerve fibre layer (RNFL) and GC complex, enabling early identification of nerve fibre loss. In contrast, FP offers a top-down view of the optic nerve head, allowing for disc cupping evaluation and broader topographic interpretation [80]. Mehta et al [81] demonstrated that combining OCT and FP could improve accuracy in detecting structural glaucomatous damage through CDR-based assessment.In the last decade, DL has significantly reshaped the landscape of glaucoma diagnostics. Early ML models often relied on handcrafted features like texture and colour histograms to classify glaucoma cases [82]. These models provided a baseline but lacked scalability and robustness in real-world scenarios.
With the rise of CNNS, researchers shifted to end-to-end learning pipelines capable of automatically extracting hierarchical visual features[10]. Li et al. [83] developed a CNN-based glaucoma detection model trained on FP images, reporting performance metrics on par with experienced ophthalmologists. Thakur et al.[84] developed a multimodal deep learning framework that combined OCT and visual field (VF) data, effectively capturing the disease’s structural and functional aspects. Similarly, Asaoka et al.[85]built on this research by employing deep regression models to forecast future visual field deterioration using only baseline OCT scans, providing a useful clinical resource for long-term monitoring.
More recently, attention-based architectures, specifically ViTs, have shown promise in retinal image analysis. Unlike CNNS, which rely on localised filters, transformer models can capture long-range dependencies and global context, making them suitable for capturing subtle glaucomatous patterns across the optic disc and RNFL regions. Liu et al.[86] applied transformer models to fundus images and reported superior classification performance and generalizability across datasets compared to conventional CNN-based baselines.
Age-Related Macular Degeneration: Age-Related Macular Degeneration is a name that suggests that it can affect the center vision part of the eye, the macula. Eyesight declines with the passage of time and can impact daily routine activities. It is complicated in the early and later stages, as the early stage shows minimal symptoms and can cause vision loss in the former stage[87]. In the case of unattended patients, AMD cases will be projected to be 288 million patients by 2040 [88]. Due to this, it poses an important public health challenge. Different case studies have used an AI-driven retinal diagnostic model to address this ailment. Mark and Andres designed the study [89],[90]to quantify and detect the abnormality of the macula in Fundus images using ML. This model helped to assess the risk of progression, along with the prediagnosis of AMD.
A comprehensive study [91] which leverages the ML algorithm, categorises the AMD stages along with the data-driven technique and develops a Computer-Aided Diagnosis framework. This model is classified into No, Intermediate, Dry, and Wet AMD, and the results are compared with those of the other studies. This model has a better result. Likewise, unlike typical ML, a deep learning algorithm is employed in different AMD patient case studies. Due to their hierarchical, deep structure, the DL model gets an improved result by interacting directly with the hidden parameters and acting as an autonomous model.
[92] presents a DL-AI-driven solution for automated AMD detection and differential diagnosis, leveraging vision transformers, data augmentation and Swin transformer through the fundus images to enhance accuracy, efficiency, and cost-effectiveness. Likewise, a DL model integrates with the local outlier Factor algorithm’s goal to classify the AMD from the OCT images. The generalizations are tested on the unseen dataset, a Duke dataset, which performs very well in the early detection of AMD[93].
Diabetic Retinopathy: When insulin production is not enough in the human body, it can raise the glucose level in the blood and lead to diabetes. Type 1 and 2 both have a high socioeconomic effect. However, type 2 requires more attention as it leads to other chronic diseases associated with heart, kidney, nose, nerve and eye diseases [94].
DR is one of the complications directly related to the eye. The complication involves swelling of the vasculature of the retina, causing the leaking of blood and fluids, and, in the later stage, it leads to vision loss if unattended. The pathophysiology of the DR risk is distributed like this: the retinal vascular leakage causes fluid accumulation in the vitreous humour. Similarly, MAs and HMs are red lesions by the same soft and hard EX. A bright lesion disrupts normal vision, and the lesion advances in the shape of Neovascularization, leading to severe vision as portrayed in the Figure 4. The commonly occurring retinal lesions include MAs, intraretinal HMs, and venous beading (characterised by alternating areas of venous dilation and constriction). Additionally, intraretinal microvascular abnormalities, hard EX (lipid deposits), and retinal neovascularization are also recognised as common types of lesions. Since these DR risks are distributed into binary and multiple classes, in multiple classes, there are often five classes, which are No, mild, moderate, severe and PDR. In contrast, in the binary, these classes are non-proliferative DR (NPDR), further subdivided into the mild, moderate, and Severe DR. Subsequently, the remaining stages are considered as PDR [95, 96, 97].
Over the past two decades, much research has been conducted to prevent DR from the classical algorithms to the modern ones [98]. A tetragonal local octa pattern method is proposed by [99]to represent the features of Fundus images. Subsequently, authors in [100] used traditional ML algorithms such as SVM, decision tree and RF. likewise, the study conducted at [101]employed the approach of Gabor wavelet encased with the Adaboost classifier to classify the DR. For the binary classification of a DR, a range of studies [102, 103, 104, 105] has been available using the CNN model, which has gained a significant result.
Furthermore, the binary and multi-classification of DR leveraging the DL model is evidenced in these studies. For instance, in [106], Rego employed InceptionV3 to detect DR using RGB and texture features. Pamadi et al. [107] developed binomial and multinomial classification models for fundus images using MobileNetV2, while Saranya et al. [108] utilised the DenseNet-121 model to identify DR from fundus images. Additionally, various architectures of EfficientNet were explored in [109]. Some studies proposed hybrid models that combined DL models for feature extraction with traditional ML models for classification. For example, Boral and Thorat [110]used InceptionV3 for representation learning in conjunction with an SVM for DR classification.
Ensemble models (EM) are widely recognised in medical imaging as an effective strategy to improve classification accuracy and model robustness by aggregating outputs from multiple base learners [111]. Jiang et al. [112]introduced an EM for DR detection, incorporating three CNN models. Furthermore, Adaboost was employed to efficiently integrate the outputs of these DL models through learned weights. The image preprocessing steps included resizing into 520x520x3, rotation, translation, mirroring, contrast adjustment, sharpness enhancement, and brightness modification.
Advancements in DL models have led to the incorporation of a self-attention mechanism, significantly enhancing both textual and visual applications and developing Vision Transformers. ViTs demonstrate superior capability in encoding spatial equivariance compared to the convolutional layers found in traditional CNNs and continue to perform exceptionally well[48, 113].
Kumar et al. [114] conducted a comparison of three DL architectures for DR detection: Transformer-based networks (specifically Swin-Transformer and Vision-Transformer), CNNs (including EfficientNet and ResNet), and a multi-layered perceptron (MLP Mixer). Their research revealed that transformer-based models outperformed the other architectures in terms of accuracy.
In another study, Yu et al., [49] introduced a Multiple Instances Learning Vision Transformer (MIL-ViT) DL model, which undergoes training in two phases: initially, it is pre-trained on a large fundus image dataset, followed by fine-tuning for the specific task of DR detection. Performance assessments utilising the [75] and RFMiD2020 [115] datasets indicated that MIL-ViT surpassed CNN models.
Adak et al. [116] developed an ensemble of transformer models that integrates four distinct models to evaluate the severity of DR. Furthermore, Gu et al. [117] proposed a DR grading model that combines a vision transformer with residual attention. This model features two primary components: (1) a feature extraction block based on the transformer, which emphasises retinal haemorrhage and exudate areas, and (2) a grading prediction block employing residual attention to capture different spatial regions corresponding to various classes. For further insights into the application of transformers in medical imaging, readers are encouraged to explore recent studies [118], [119].
These diseases and their detection are pivotal and are used in the vivio and vito, and have established a foundation for the field of oculomics. However, diseases covered through the above data sets may not have the potential to get deep insight into the underlying mechanism of genotype-wide association. These are primarily focused on phenotypic analysis, which limits their utility for genotypic studies. To fully realise the prospective of oculomics, the acquisition of datasets needs to be extended to include cohort-based studies that integrate clinical patient records, high-throughput data, and structural and functional imaging data. Such comprehensive datasets would enable researchers to explore the genetic and molecular underpinnings of ocular and systemic diseases, paving the way for personalised medicine.

### v2

@@ II-B8 Overview on the Evolution of AI
In a nutshell, the emergence of AI has reshaped the dynamics of almost every field, specifically healthcare. The short description, limitations, evolution, and relevance to medicine are shown in Figure 4, and some related stuff is discussed here. In 1956, the emergence of AI took its main start from the Dartmouth College conference, which mainly consisted of symbolic AI, in other words, a rule-based system. However, due to the computational limit, the transition of the rule-based system into the classical ML occurred from 1966 to 1970. The wave of backpropagation was another milestone in 1980; despite the emergence of this technique, AI was not getting full attention due to the funding constraints, along with the expected results Perez-Lopez et al. (2024). The era 1980-1990 is considered a winter era for AI, and again, the domain is shifted to statistical and classical ML. The Chronology of LENET Architecture, which spans from 1990 to 2012, laid the foundation for the emergence of the DL model and paved a new horizon for the future of AI LeCun et al. (2015); Rawat & Wang (2017). The DL model is a seismic change that is now non-stop in the current era, and day-to-day, new algorithms and models are emerging. The transition of AI from a theoretical to a practical tool has changed the dynamics of the research, academia and industrial world.
The capabilities of AI now extend from the basic models to the complex, whether supervised or unsupervised, transforming research, clinical practice, and industry alike. Within ophthalmology, this evolution intersects directly with the trajectory of retinal imaging. While classical imaging techniques laid the foundation for retinal diagnostics, they were largely limited to descriptive or handcrafted analyses. These approaches successfully characterised ocular lesions and vascular features but lacked scalability, sensitivity, and integration with systemic health. The rise of artificial intelligence has transformed this landscape, enabling automated segmentation, feature extraction, and predictive modelling at a scale unattainable with manual methods. Section III explores how modern AI-driven methods build upon these classical foundations, translating retinal images into quantitative biomarkers with applications that extend beyond ocular disease to systemic health.
III Modern Techniques: The Rise of Oculomics
As mentioned earlier, the eye disease is retained in the retina. The same is also connected and affects the important parts of the body, such as the brain, heart, kidneys, etc. This type of disease is called systemic disease, a condition that affects the entire body or several organs, not just one part. The linking of eyes with systemic diseases is coined oculomics Zhou et al. (2022); Wagner et al. (2020). The potential behind oculomics has expanded further by leveraging the existing retinal imaging research with the advancement of multiple data modalities. This can be gained by revealing the medical modalities. The modalities used so far in the field of ophthalmology are OCT, fluorescence angiography (FA), and CFP to reveal structural and functional changes in the ocular system Ringel et al. (2021); Honavar (2022). These modalities help provide enormous retinal imaging data and are widely used in the real world of medical vision. These imaging data emerged as the dataset used to mitigate the burden of eye diseases with respect to monitoring, therapeutic, prognostic, and diagnostic, respectively Ringel et al. (2021). The datasets are used in different case studies individually and combined for monitoring and identifying a variety of eye diseases, such as DR, glaucoma, macular edema, and cataracts, etc. The different retinal datasets are publicly available to detect and classify eye diseases on time. These datasets have been very helpful so far and are used for different objectives of classical ocular diseases.
Figure 2: The schematic illustrates the spectrum of retinal imaging technologies. Modalities shown on a pink background (CFP, OCT, FA, OCTA) represent the core, widely adopted clinical tools that underpin current diagnostic practice. Modalities grouped on a white background (WFDI, HFI, AO, AO-OCT, AO-SLO, cSLO) reflect advanced or emerging approaches that extend the field of view, resolve cellular or biochemical detail, or enhance functional assessment. While these latter techniques remain largely research-focused, they point to the expanding potential of multimodal imaging for linking retinal structure with systemic disease processes.
III-A Retinal Imaging Techniques
Retinal imaging technologies have evolved rapidly, trans- forming our ability to capture both structure and function of the eye at increasing resolution. Among the classical modalities, colour fundus photography remains one of the most widely applied tools in both clinical practice and large-scale epidemiological studies. Its strengths lie in simplicity, low cost, and broad availability, making it indispensable for screening conditions such as diabetic retinopathy, hypertensive retinopathy, and retinal vein occlusion. For decades, CFP has served as the standard tool in population-based studies, enabling quantification of vascular parameters such as caliber, tortuosity, fractal dimension, width, and focal narrowing. The rationale is that the retina, as the only part of the central nervous system visible non-invasively, provides direct access to microvascular and neural health Tan et al. (2022). Modern image analysis platforms now extend their utility: SIVA (Singapore I Vessel Assessment), IVAN (Integrative Vessel Analysis), and VAMPIRE (Vascular Assessment and Measurement Platform for Images of the Retina) offer semi-automated measurements of vascular geometry Mautuit et al. (2022), while AutoMorph provides a fully automated, end-to-end pipeline for vessel segmentation and trait extraction Zhou et al. (2022). The major limitation of CFP remains its two-dimensional nature, which lacks depth resolution and reduces sensitivity to subtle intraretinal changes.
Optical coherence tomography introduced a step change in retinal diagnostics by allowing cross-sectional, micrometer-resolution imaging of the retinal layers. The progression from time-domain to spectral-domain OCT has greatly improved speed, resolution, and reproducibility, supporting the quantitative evaluation of the retinal sublayers Geevarghese et al. (2021); Pazos et al. (2021). OCT is now the gold standard for detecting macular edema, glaucoma-related thinning, and AMD, although its cost and sensitivity to motion artifacts remain challenges. The functional extension, OCT angiography (OCTA), non-invasively visualizes retinal and choroidal microvasculature by detecting red blood cell motion. OCTA generates depth-resolved vascular maps and vessel density metrics that are particularly valuable in ischemic conditions and neovascular disease Greig et al. (2020). Yet, OCTA can miss slow or turbulent flow, is prone to segmentation errors, and typically covers a narrower field than wide-field angiography.
Fluorescein angiography has long been regarded as a core imaging tool for assessing retinal vasculature. By injecting sodium fluorescein dye, FA highlights vascular leakage, non-perfusion, and neovascularization, making it essential in diagnosing diabetic macular edema, retinal vein occlusion, and neovascular AMD Ricardi et al. (2024). Despite its diagnostic utility, FA is invasive, time-consuming, and associated with risks such as nausea and, rarely, anaphylaxis, prompting the development of non-invasive alternatives like OCTA.
Beyond these established modalities, a number of emerging technologies extend the diagnostic horizon. Wide-field digital imaging (WFDI) captures up to 200° Witmer & Kiss (2013) of the retina in a single shot, proving especially useful in identifying peripheral changes in diabetic retinopathy and for neonatal screening Desurmont et al. (2023). Hyperspectral imaging (HFI) provides biochemical and metabolic information by measuring reflectance at multiple wavelengths, revealing signals such as hemoglobin oxygenation and amyloid deposits associated with Alzheimer’s disease Hadoux et al. (2019). These tools remain largely research-focused but hold promise for systemic disease characterization.
Adaptive optics (AO) should be understood as an enabling technology rather than a standalone clinical modality. By correcting ocular aberrations, AO achieves near-cellular resolution, and when coupled with imaging platforms such as scanning laser ophthalmoscopy (AO-SLO) or OCT (AO-OCT), it allows direct visualization of photoreceptors, retinal pigment epithelial cells, and fine capillaries in vivo. AO has provided unprecedented insights into inherited retinal diseases, including Stargardt disease, retinitis pigmentosa Georgiou et al. (2018), and cone–rod dystrophy Wong (2004), and also shows potential for functional imaging by enabling the tracking of erythrocyte movement and local blood flow dynamics at the capillary level Tan et al. (2022). However, its technical complexity and limited commercial availability currently confine AO to research laboratories rather than routine clinics. In parallel, confocal scanning laser ophthalmoscopy (cSLO) has enhanced classical retinal imaging, supporting modalities such as FA and fundus autofluorescence, with recent refinements including multicolor and ultra-widefield capabilities Horie et al. (2021). AO and cSLO extend imaging into both cellular and functional domains, complementing OCT and CFP within multimodal frameworks.
Today, the landscape of retinal imaging ranges from classical modalities such as CFP, FA, and OCT, which remain central to clinical care, to newer innovations including OCTA, WFDI, HFI, AO-enabled systems, and cSLO-based platforms that expand the scope of investigation into vascular, biochemical, cellular, and functional domains. As illustrated in Figure 2, this spectrum can be organised into core modalities that underpin current diagnostic practice and emerging approaches that extend capabilities for research and multimodal integration. Each approach contributes unique strengths but also carries limitations, underscoring why multimodal combinations are increasingly applied. These limitations may be further compounded in patients with unstable fixation or ocular motility disorders such as nystagmus, where motion artefacts can compromise image quality and reduce the reliability of quantitative outcomes. Nevertheless, the convergence of classical and emerging technologies has generated an unprecedented volume of high-resolution ocular data, enabling the creation of large-scale repositories. These resources now play a central role in validating AI-driven pipelines and advancing oculomics toward clinical translation, as explored in the next section on retinal datasets.
III-B Publicly Available Retinal Datasets
In the present day, the data is used as a tool for competing, from small tech companies to giants and from developing to developed countries. Data is valuable and works as a gold mine in the epoch of GEN-AI. In this context, the available data is getting more attention if found in large amounts, specifically in the field of the medical sciences. This data can emerge as a dataset which can be used further in the real world of health research and drive innovations and discoveries Khan, Liu, Nath, Korot, Faes, Wagner, Keane, Sebire, Burton & Denniston (2021). Furthermore, gold-standard imaging techniques provide the opportunity for the creation of a dataset used as a benchmark for the validation and testing of real clinical trials in situ and in vivo. The advancement of retinal imaging techniques has created numerous publicly available datasets, which have become indispensable resources for researchers and physicians. These datasets are widely used for lesion detection, disease classification, and vessel segmentation Li et al. (2021). For example, datasets such as STARE Hoover et al. (2000),DRIVE Staal et al. (2004),DIARETDB1 Kauppi et al. (2007) ,CHASE DB1, Owen et al. (2009) ,HRF Budai et al. (2013), E-ophtha Decenciere et al. (2013) , MESSIDOR, MESSIDOR-2 Decencière et al. (2014) ,KAGGLE-EyePACS Dugas et al. (2015) ,IDRiD Porwal et al. (2018), KAGGLE-APTOS Karthik et al. (2019) , and DDR Li et al. (2019) have been instrumental in developing algorithms for detecting and classifying diabetic retinopathy, segmenting retinal vessels, and identifying other ocular pathologies.
In this survey, therefore, we transition from the dataset overview into a focused discussion of key localized diseases: beginning with glaucoma, where structural deterioration of the optic nerve is a primary concern; followed by age-related macular degeneration, which involves degeneration of the macula; and finally diabetic retinopathy, where vascular abnormalities and lesion characteristics dominate. Each disease presents unique imaging biomarkers and machine learning challenges, necessitating tailored approaches in algorithm design and data representation.
Figure 3: An annotated fundus image reproduced from Li et al. (2021), showing pathological markers including microaneurysm, haemorrhage, exudates, and anatomical regions like the fovea and optic disc.
Glaucoma: Glaucoma is a long-term and gradually worsening condition of the optic nerve that results in permanent loss of vision and stands as one of the primary causes of blindness worldwide. It affects an estimated 76 million people as of 2020 and is projected to impact over 111.8 million by 2040 Tham et al. (2014). Elevated intraocular pressure is often linked to the disease, but many patients also experience normal-tension glaucoma, illustrating the condition’s complex origins. The disease affects both the structural and functional health of the eye, especially through the degeneration of retinal ganglion cells (GC) and harm to the optic nerve head. Clinically, a key biomarker for diagnosing glaucoma is the cup-to-disc ratio, which quantifies the size of the optic cup relative to the optic disc and reflects the cupping of the optic nerve Wagner et al. (2022); Sahu (2024).
Two primary imaging modalities, CFP and OCT, are widely used to detect and monitor glaucomatous changes. OCT provides high-resolution cross-sectional views of the RNFL and GC complex, enabling the early identification of nerve fiber loss. In contrast, CFP offers a top-down view of the optic nerve head, allowing for disc cupping evaluation and broader topographic interpretation Coan et al. (2023). Mehta et al. (2021) demonstrated that the combination of OCT and CFP scans could improve accuracy in detecting structural glaucomatous damage by assessing the cup-to-disc ratio. In the last decade, deep learning has significantly reshaped the landscape of glaucoma diagnostics. Early ML models often relied on handcrafted features like texture and colour histograms to classify glaucoma cases Bock et al. (2010). These models provided a baseline but lacked scalability and robustness in real-world scenarios.
With the rise of CNNS, researchers shifted to end-to-end learning pipelines capable of automatically extracting hierarchical visual featuresTing et al. (2019). In a study, Li et al. (2018) developed a CNN-based glaucoma detection model trained on fundus images, reporting performance metrics on par with experienced ophthalmologists. The research conducted by Kihara et al. (2022) established a multimodal deep learning framework that integrated Optical Coherence Tomography and visual field data. This approach effectively captured both the structural and functional aspects of the disease. Similarly, Xu et al. (2021) built on this research by employing deep regression models to forecast future visual field deterioration using only baseline OCT scans, providing a useful clinical resource for long-term monitoring.
More recently, attention-based architectures, specifically ViTs, have shown promise in retinal image analysis. Unlike CNNS, which rely on localised filters, transformer models can capture long-range dependencies and global context, making them suitable for capturing subtle glaucomatous patterns across the optic disc and RNFL regions. Playout et al. (2022) applied transformer models to fundus images and reported superior classification performance and generalizability across datasets compared to conventional CNN-based baselines.
Age-Related Macular Degeneration: Age-Related Macular Degeneration is a name that suggests that it can affect the center vision part of the eye, the macula. Eyesight declines with the passage of time and can impact daily routine activities. It is complicated in the early and later stages, as the early stage shows minimal symptoms and can cause vision loss in the former stage Gao et al. (2024). In the case of unattended patients, AMD cases will be projected to be 288 million patients by 2040 Wong et al. (2014). Due to this, it poses an important public health challenge. Different case studies have used an AI-driven retinal diagnostic model to address this ailment. Mark and Andres designed the study van Grinsven et al. (2013),García-Floriano et al. (2019) to quantify and detect the abnormality of the macula in Fundus images using ML. This model helped to assess the risk of progression, along with the prediagnosis of AMD.
A comprehensive study Abd El-Khalek et al. (2024) which leverages the ML algorithm, categorises the age-related macular degeneration stages along with the data-driven technique and develops a Computer-Aided Diagnosis framework. This model is classified into No, Intermediate, Dry, and Wet AMD, and the results are compared with those of the other studies. This model has a better result. Likewise, unlike typical machine learning, a deep learning algorithm is employed in different case studies of AMD patients. Due to their hierarchical, deep structure, the DL model gets an improved result by interacting directly with the hidden parameters and acting as an autonomous model.
A deep learning-AI-driven solution has been developed for automated AMD detection and differential diagnosis, leveraging vision transformers, data augmentation, and Swin transformers to improve accuracy, efficiency, and cost-effectiveness Xu et al. (2023). Likewise, a deep learning model integrates with the local outlier Factor algorithm’s goal to classify the age-related macular degeneration from the OCT scans. The generalizations are tested on the unseen dataset, a Duke dataset, which performs very well in the early detection of AMD He et al. (2022).
Diabetic Retinopathy: When insulin production is not enough in the human body, it can raise the glucose level in the blood and lead to diabetes. Type 1 and 2 both have a high socioeconomic effect. However, type 2 requires more attention as it leads to other chronic diseases associated with heart, kidney, nose, nerve and eye diseases Azeem et al. (2022).
DR is one of the complications directly related to the eye. The complication involves swelling of the retinal vasculature, causing blood and fluid to leak, and, at a later stage, it leads to vision loss if left unattended. The pathophysiology of the DR risk is distributed like this: the retinal vascular leakage causes fluid accumulation in the vitreous humour. Similarly, MA and HM are red lesions, and by the same soft and hard EX, a bright lesion disrupts normal vision, and the lesion advances in the shape of Neovascularization, leading to severe vision as portrayed in Figure 3 Wei et al. (2022). The commonly occurring retinal lesions include MAs, intraretinal HMs, and venous beading (characterised by alternating areas of venous dilation and constriction). Additionally, intraretinal microvascular abnormalities, hard EXs (lipid deposits), and retinal neovascularization are also recognised as common types of lesions Li et al. (2021). Since these DR risks are distributed into binary and multiple classes, in multiple classes, there are often five classes, which are No, mild, moderate, severe and PDR Sarki et al. (2020). In contrast, in the binary, these classes are non-proliferative DR (NPDR), further subdivided into the mild, moderate, and Severe DR. Subsequently, the remaining stages are considered as PDR Khan, Khan, Khan, Rehman, Shah, Qummar, Ali & Pack (2021).
Over the past two decades, much research has been conducted to prevent DR from the classical algorithms to the modern ones Yang et al. (2022). A tetragonal local octa pattern method was proposed to represent features of fundus images Nazir et al. (2019). Building on this, Gayathri et al. (2021)applied traditional ML algorithms such as SVM, decision trees, and random forests for classification. Likewise, the study conducted at Washburn et al. (2020) employed the approach of Gabor wavelet encased with the Adaboost classifier to classify the DR. For the binary classification of a DR, a range of studies Xu et al. (2017); Esfahani et al. (2018); Zago et al. (2020); Suriyal et al. (2018) has been available using the CNN model, which has gained a significant result.
Furthermore, the binary and multi-classification of DR leveraging the DL model is evidenced in these studies. For instance, in Rêgo et al. (2021), InceptionV3 was employed to detect DR using RGB and texture features. Pamadi et al. (2022) developed binomial and multinomial classification models for fundus images using MobileNetV2, while Saranya et al. (2022) utilised the DenseNet-121 model to identify DR from fundus images. Additionally, various architectures of EfficientNet were explored in Mudaser et al. (2021). Some studies proposed hybrid models that combined DL models for feature extraction with traditional ML models for classification. For example, Boral & Thorat (2021)used InceptionV3 for representation learning in conjunction with an SVM for DR classification.
Ensemble models (EMs) have long been recognised in medical image analysis as an effective way to improve classification accuracy and robustness by combining the outputs of multiple base learners Hassan et al. (2024). In the case of DR detection, Jiang et al. (2019) demonstrated this approach with an ensemble of three CNNs, where AdaBoost was used to optimise the integrated predictions. More recently, attention has shifted toward transformer-based architectures, which employ self-attention mechanisms to capture long-range spatial relationships that CNNs often miss. The introduction of Vision Transformers Dosovitskiy et al. (2021) marked a turning point, and subsequent studies have highlighted their value in medical imaging applications Amin et al. (2025). Comparative evaluations confirm this advantage: Kumar & Karthikeyan (2021) found that Swin-Transformer and Vision-Transformer consistently outperformed CNN-based models such as EfficientNet and ResNet, as well as MLP-Mixer architectures. Building on this progress, Yu et al. (2021) introduced the Multiple Instance Learning Vision Transformer (MIL-ViT), which was pretrained on large fundus datasets before fine-tuning for DR tasks. When tested in APTOS2019 Karthik et al. (2019) and RFMiD2020 Pachade et al. (2020), MIL-ViT achieved superior accuracy compared to conventional CNNs.
An ensemble of transformer models was developed that integrates four distinct networks to evaluate the severity of DR Adak et al. (2023). Furthermore, Gu et al. (2023) proposed a DR grading model that combines a vision transformer with residual attention. This model features two primary components: (1) a feature extraction block based on the transformer, which emphasises retinal haemorrhage and exudate areas, and (2) a grading prediction block employing residual attention to capture different spatial regions corresponding to various classes. For further insights into the application of transformers in medical imaging, readers are encouraged to explore recent studies Parvaiz et al. (2023), Ye et al. (2023).
These diseases and their detection are pivotal and are used in vivo and in vitro, and have established a foundation for the field of oculomics. However, diseases covered in the above data sets may not have the potential to provide deep insight into the underlying mechanism of genotype-wide association. These are primarily focused on phenotypic analysis, which limits their utility for genotypic studies. To fully realise the prospective of oculomics, the acquisition of datasets needs to be extended to include cohort-based studies that integrate clinical patient records, high-throughput data, and structural and functional imaging data. Such comprehensive datasets would enable researchers to explore the genetic and molecular underpinnings of ocular and systemic diseases, paving the way for personalised medicine. Importantly, one of the most promising directions for such integration lies in the automated quantification of retinal vascular morphology, as the retinal circulation directly reflects systemic vascular health and provides measurable traits that can bridge imaging, genetics, and multi-omics frameworks